# Physical Processes Behind the Co-Evolution of Halos, Galaxies and Supermassive Black Holes in the IllustrisTNG Simulation

### v2

@@ 4.3.1 The self-regulated growth
Figure 7: The relation between the power-law index, $k$, of the $M_{\rm BH}$-$M_{*}$ history during Phase-3 and the ratio, $M_{\rm BH}/M_{*}$, at $t_{23}$, the beginning of Phase-3. Dots in the bottom left panel represent individual galaxies, color-coded by the value of $M_{\rm BH}$ at $t_{23}$. The Spearman correlation coefficient and the $p$-value are indicated in the label. Black dashed line shows a linear fit, which gives a slope of about $-3.0$. Top left and bottom right panels show the two marginal distributions, respectively. See §4.3.1 for the details of how this figure implies a self-regulated growth.
The self-regularization nature can be seen from Eq. (14) by considering a perturbation on the gas properties and the consequence of such perturbation on the feedback energy ($E_{\rm agn}$) and the binding energy ($E_{\rm bind}$). If gas inflow is temporarily enhanced by, e.g. the fluctuations of halo assembly or the disk instability, Bondi accretion will be accelerated, leading to a trajectory in the $M_{\rm BH}$-$M_{*}$ plane that resembles the one in Phase-2 and producing an over-massive SMBH. The feedback output then rises, leading to an inequality, $E_{\rm agn}>\eta_{\rm eff}^{-1}E_{\rm bind}$. The gas surrounding the SMBH will then be evacuated and the growth of SMBH will be stalled. Inversely, if the gas inflow is suppressed by, e.g. the temporary reduction of halo accretion or the consumption of gas by star formation, Bondi accretion will be stopped, leading to a shallower trajectory in the $M_{\rm BH}$-$M_{*}$ plane. The inequality will be inverted as $E_{\rm agn}<\eta_{\rm eff}^{-1}E_{\rm bind}$, and so the gas will be accumulated again, and the growth of the SMBH will be accelerated. The discussion here implies that the evolution of the galaxy and its SMBH closely follows a trajectory defined by the equilibrium equation $E_{\rm agn}=\eta_{\rm eff}^{-1}E_{\rm bind}$.
To convert this prediction into an explicit relation, we use the virial theorem, and approximate the binding energy as $E_{\rm bind}\sim M_{\rm g}V_{\rm g}^{2}$. Combined with the feedback energy given by Eq. (4.2), we obtain
@@
where the scaling factor $f\equiv\eta_{\rm eff}^{-1}\eta_{\rm thermal}^{-1}f_{\rm gas}f_{*}^{-5/3}$. The values of $\eta_{\rm eff}$ and $\eta_{\rm thermal}$ depend on the implementation of the feedback model in the simulation, as described in §4.2. One interesting feature of $f_{*}$, as suggested by observations (e.g. Yang et al., 2007, 2012; Behroozi et al., 2019), is the flattening at the critical mass, $M_{\rm h,c}\approx 10^{12}\,{\rm M_{\odot}}$ (or $M_{\rm*,c}\approx 10^{10.5}\,{\rm M_{\odot}}$), where $f_{*}$ appears to be independent of halo mass to the first order. The gas fraction, $f_{\rm gas}$, is also expected to peak at this critical mass, as suggested by (Terrazas et al., 2020; Voit et al., 2024). Thus, at around $M_{\rm*,c}$, all these coefficients are roughly constant, with only moderate redshift dependence introduced by the relation between $M_{\rm h}$ and $V_{\rm h}$. Consequently, the equilibrium equation predicts a power-law relation, $M_{\rm BH}\propto M_{*}^{5/3}$. At lower or higher mass scales, $f_{*}$ and $f_{\rm gas}$ cannot be viewed as a constant, and the power-law index may deviate from $5/3$. Note that the prediction of the scaling relation here is consistent with the semi-analytical model of Mo et al. (2024) and the observationally calibrated toy model of Hong et al. (2023), although they focus on the formation of bulge component and quenched galaxies, respectively.
@@
The example shown by the red dashed curve in Fig. 2 demonstrates the regulation over the equilibrium scaling relation. At $t_{\rm 23}$, the SMBH mass is slightly overshot relative to the power-law scaling with index $\gamma=5/3$. During Phase-3, the galaxy is pulled back towards the scaling, and finally comes to a region in the $M_{\rm BH}$-$M_{*}$ plane where other galaxies (shown by grey curves) converge. To statistically quantify the self-regulation nature, in Fig. 7, we show $k$, the power-law index of the $M_{\rm BH}$-$M_{*}$ relation during Phase-3 obtained from the best-fit function (Eq. 2), and $M_{\rm BH}(t_{23})/M_{*}(t_{23})$, the ratio of SMBH mass and stellar mass at the beginning of Phase-3, for individual galaxies, as well as the marginal distributions. The distribution of $M_{\rm BH}(t_{23})/M_{*}(t_{23})$ has a median of $\approx 10^{-2.4}$, approaching the typical value of galaxies at $z=0$. The distribution of $k$ has a median of $\sim 1.35$, smaller than the value of $5/3$ that Eq. (4.3.1) would imply if both $f_{\rm gas}$ and $f_{*}$ are constant. A strong anti-correlation between $k$ and $M_{\rm BH}(t_{23})/M_{*}(t_{23})$ is seen from the scattered points, and confirmed by the negative (rank-based) Spearman correlation coefficient. The dashed line in Fig. 7 shows a linear regression of $k$ on $\log\,\left[M_{\rm BH}(t_{23})/M_{*}(t_{23})\right]$, and has a slope of $\approx-3.0$. This slope indicates that, if a SMBH deviates from the equilibrium scaling relation (Eq. 4.3.1), it is pulled back once $M_{*}$ grows for about $0.3\,{\rm dex}$. For a typical main-sequence galaxy at $z\approx 0$ with sSFR $\sim 1\,{\rm Gyr}^{-1}$, this pull-back time scale is $\sim 2\,{\rm Gyr}$, shorter than the Phase-3 duration of most galaxies (Fig. 2). Hence, most SMBHs are well self-regulated during Phase-3, which erases the footprints of their early growth, including the seeding process. This explains the tight relation between $M_{\rm BH}$ and $M_{*}$ at $z=0$, as shown by the black solid curve in Fig. 2. Observations also suggest a relation between $M_{\rm BH}$ and $M_{*}$ (or $M_{\rm*,bulge}$) in the local Universe (e.g. Kormendy & Ho, 2013; Greene et al., 2020; Zhang et al., 2024), which we will discuss in §5.1.
@@
The similar trend of sHAR, sSFR, and sBHAR during Phase-3 shown in the top left panel of Fig. 6 suggests that the rate of evolution of a galaxy in the $M_{\rm BH}$-$M_{\rm*}$ plane is controlled by the accretion rate of the host halo. The regulation by the AGN feedback bridges the growth of SMBH with its environment, transfers the growth pattern of the host halo to the galaxy and SMBH, and leads to the establishment of the scaling relation derived in Eq. (4.3.1). The sBHAR decreases with time during Phase-3, as it follows the sHAR that in general decreases with time. Once a certain mass threshold is reached, the mode of AGN feedback implemented in TNG makes a transition, as we will discuss in the following sections.
@@ 4.3.2 The trigger of the kinetic-mode feedback
Figure 8: Distribution of the properties relevant to the quenching of galaxies. Top panels show the distributions of individual galaxies in the $t_{\rm kin}$-$t_{\rm vmax,peak}$ (left) and $t_{\rm q}$-$t_{\rm vmax,\,peak}$ left planes, color-coded by their $M_{\rm h}$ at $z=0$. Bottom panels show the histograms of $t_{\rm q}-t_{\rm kin}$, the time lag of quenching after the trigger of the kinetic-mode feedback. The grey histogram in each panel shows the result for a subsample of galaxies in a given range of $M_{\rm h}$ at $z=0$. This subsample is further divided into three bins by $f^{[t_{\rm kin},t_{\rm q}]}_{\rm*,ex}$, the ex-situ growth of stellar mass between $t_{\rm kin}$ and $t_{\rm q}$, and their histograms are shown in different colors. The number of galaxies in each bin is labeled in the legend. See §4.3.2 and §4.3.3 for the definitions of these times and their implications for galaxy quenching.
@@
$\chi_{\rm thres}\equiv\min\left[0.02\left(\frac{M_{\rm BH}}{10^{8}\,{\rm M_{\odot}}}\right)^{2},0.1\right]\,,$ (17)
@@
For low-mass SMBHs ($M_{\rm BH}\lesssim 10^{8}\,{\rm M_{\odot}}$), the threshold $\chi_{\rm thres}$ is close to zero, and thus the transition occurs only if the sHAR approaches zero. A universal feature of halo assembly is the presence of two phases: a fast-accretion phase at high redshift, when the halo potential rapidly deepens, and a slow-accretion phase at low redshift, when the halo potential remains constant or even declines (e.g. Zhao et al., 2003). Thus, for the SMBH in a low-mass halo to trigger the kinetic-mode feedback, its host halo must transit to the slow phase. To verify that the halo transition indeed leads to the trigger of the kinetic-mode feedback, we follow Zhao et al. (2003) to define the halo transition time as $t_{\rm vmax,\,peak}$, the time when the maximum of the circular velocity, $V_{\rm max}$, reaches the peak, and we compare it to $t_{\rm kin}$, the time when the kinetic-mode feedback is triggered. The results are shown in the top left panel of Fig. 8 for individual halos, color-coded according to their halo masses at $z=0$. For halos with $M_{{\rm h},z=0}\lesssim 10^{12.5}\,{\rm M_{\odot}}$, $t_{\rm vmax,\,peak}$ correlates well with $t_{\rm kin}$, indicating that the transition of the AGN feedback is indeed driven by the transition of halo accretion.
For massive SMBHs ($M_{{\rm BH},z=0}\gtrsim 7\times 10^{8}\,{\rm M_{\odot}}$, or $M_{{\rm h},z=0}\gtrsim 10^{13}\,{\rm M_{\odot}}$), the threshold $\chi_{\rm thres}$ approaches the upper limit $0.1$. The transition of these halos can occur even when they are fast accreting. To derive the condition for their transition, we follow Dekel et al. (2023, see their §5.1) to approximate the specific accretion rate of a fast-accreting halo as
@@ 4.3.3 The quenching of galaxies
Figure 9: The sufficient condition for the quenching of galaxies. Left panel shows the distribution of galaxies in the $\dot{E}_{\rm cool,sub}$-$\dot{E}_{\rm kin}$ plane, where $\dot{E}_{\rm cool,sub}$ is the cooling rate of all non-star-forming gas cells in the subhalo, and $\dot{E}_{\rm kin}$ is the energy output rate from kinetic-mode AGN. Contours with a given color include galaxies at a given time point in their evolution histories, from inner to outer, encompassing 10% – 95% of the sample. Limited by the data storage, cooling rates are only available in the 20 “full snapshots” of TNG. Thus, the contours for $t_{\rm kin}$ and $t_{\rm q}$ only include the galaxies whose cooling rate can be obtained, leading to a reduction of the sample size to 199 and 304, respectively. Dashed line indicates where these two rates are equal. Right panel shows the histograms of the ratio, $\dot{E}_{\rm kin}/\dot{E}_{\rm cool,sub}$, at different time points. Dashed lines indicate the medians of the distributions. See §4.3.3 for the details.
The feedback energy in the kinetic mode implemented by TNG affects the gas effectively by both ejection and prevention, and can quickly reduce the amount of star-forming gas in the galaxy and beyond (e.g. Zinger et al., 2020; Shi et al., 2022; Martín-Navarro et al., 2021). Consequently, galaxies are expected to be quenched quickly after the transition of the feedback mode. In the top right panel of Fig. 8, we show $t_{\rm q}$, the time when a galaxy is quenched, and compare it to $t_{\rm vmax,\,peak}$ for individual galaxies. For all galaxies, $t_{\rm q}$ is systematically later (larger) than $t_{\rm kin}$, indicating that the kinetic-mode AGN feedback is a necessary condition for quenching central galaxies, and that galaxies may sustain their star formation for some time after the trigger of the kinetic-mode feedback.
@@
where $\Delta M_{\rm*,ex}^{[t_{\rm kin},t_{\rm q}]}$ is the mass of stars formed ex-situ during the time interval between $t_{\rm kin}$ and $t_{\rm q}$ and fallen into the galaxy before $t_{\rm q}$ via either stripping or merger, and $M_{*}(t_{\rm q})$ is the stellar mass of the galaxy at $t_{\rm q}$. The distribution of the time lag of quenching for each subsample is shown in the bottom panels of Fig. 8 by a colored histogram. For galaxies with $f_{\rm*,ex}^{[t_{\rm kin},t_{\rm q}]}<5\%$, the time lag is sharply peaked within $1\,{\rm Gyr}$, regardless of the halo mass. This demonstrates that the kinetic-mode feedback is indeed effective in quenching the galaxy in the absence of merger-induced gas inflow. However, for galaxies with strong mergers ($f_{\rm*,ex}^{[t_{\rm kin},t_{\rm q}]}>30\%$), star formation can sustain a long time, even if the kinetic-mode feedback is triggered. The quenching of a galaxy is thus delayed to a time when the frequency of mergers falls below a certain level, depending on the sSFR threshold used to define the quenching (§2). A similar conclusion was reached by Terrazas et al. (2020), where the cumulative energy of AGN is computed and compared to the binding energy of the gas in a galaxy. Galaxies are found to be quenched only when the feedback energy exceeds about 100 times the binding energy of the gas, indicating that most of the feedback energy is lost due to radiative cooling.
@@
The cease of the merger marks the end of the stochastic creation of star-forming gas, leaving the cooling of hotter gas in a halo being the only channel for the central galaxy to maintain its star formation. In this case, the gas surrounding the SMBH is dominated by that in the hot phase, as shown in the bottom right panel of Fig. 6. If the energy injection by the kinetic-mode feedback remains strong even when the amount of gas available for the SMBH to accrete is small, as is the case in TNG because of its large feedback efficiency at low accretion rates (Weinberger et al., 2017, see their §2.3), the amount of the gas within the galaxy will continue to decrease due to both ejection and prevention, causing the rates of gas cooling and star formation to slow down. As shown in the top-right panel of Fig. 6, the effective sound speed of the gas during Phase-3 continues to increase, while the density decreases, both leading to a decrease in $A_{\rm s}$ and thus to the growth of the SMBH and the strength of its feedback. Such an evolution stops once the feedback of AGN is reduced to a level comparable to the cooling rate of the gas. This final stage, where cooling, star formation, SMBH growth, and AGN feedback are all slow, produces a quenched state to be consistent with observations.
@@
The above argument leads to a sufficient condition for a galaxy without significant mergers to be quenched, which is that the kinetic-mode AGN feedback balances the cooling of gas:
@@
[…]

### v3

@@ 4.3.1 The self-regulated growth
Figure 7: The relation between the power-law index, $k$, of the $_{\rm BH}-MM_{*}$ history during Phase-3 and the ratio, $M_{\rm BH}/M_{*}$, at $t_{23}$, the beginning of Phase-3. Dots in the bottom left panel represent individual galaxies, color-coded by the value of $M_{\rm BH}$ at $t_{23}$. The Spearman correlation coefficient and the $p$-value are indicated in the label. Black dashed line shows a linear fit, which gives a slope of about 3.0-3.0. Top left and bottom right panels show the two marginal distributions, respectively. See §4.3.1 for the details of how this figure implies a self-regulated growth.
The self-regularization nature can be seen from Eq. (14) by considering a perturbation on the gas properties and the consequence of such perturbation on the feedback energy ($E_{\rm agn}$) and the binding energy ($E_{\rm bind}$). If gas inflow is temporarily enhanced by, e.g. the fluctuations of halo assembly or the disk instability, Bondi accretion will be accelerated, leading to a trajectory in the $_{\rm BH}-MM_{*}$ plane that resembles the one in Phase-2 and producing an over-massive SMBH. The feedback output then rises, leading to an inequality, $E_{\rm agn}>\eta_{\rm eff}^{-1}E_{\rm bind}$. The gas surrounding the SMBH will then be evacuated and the growth of SMBH will be stalled. Inversely, if the gas inflow is suppressed by, e.g. the temporary reduction of halo accretion or the consumption of gas by star formation, Bondi accretion will be stopped, leading to a shallower trajectory in the $_{\rm BH}-MM_{*}$ plane. The inequality will be inverted as $E_{\rm agn}<\eta_{\rm eff}^{-1}E_{\rm bind}$, and so the gas will be accumulated again, and the growth of the SMBH will be accelerated. The discussion here implies that the evolution of the galaxy and its SMBH closely follows a trajectory defined by the equilibrium equation $E_{\rm agn}=\eta_{\rm eff}^{-1}E_{\rm bind}$.
To convert this prediction into an explicit relation, we use the virial theorem, and approximate the binding energy as $E_{\rm bind}\sim M_{\rm g}V_{\rm g}^{2}$ at the galaxy scale. Combined with the feedback energy given by Eq. (4.2), we obtain
@@
where the scaling factor $f\equiv\eta_{\rm eff}^{-1}\eta_{\rm thermal}^{-1}f_{\rm gas}f_{*}^{-5/3}$. The values of $\eta_{\rm eff}$ and $\eta_{\rm thermal}$ depend on the implementation of the feedback model in the simulation, as described in §4.2. One interesting feature of $f_{*}$, as suggested by observations (e.g. Yang et al., 2007, 2012; Behroozi et al., 2019), is the flattening at the critical mass, $M_{\rm h,c}\approx 10^{12}\,{\rm M_{\odot}}$ (or $M_{\rm*,c}\approx 10^{10.5}\,{\rm M_{\odot}}$), where $f_{*}$ appears to be independent of halo mass to the first order. The gas fraction, $f_{\rm gas}$, is also expected to peak at this critical mass, as suggested by (Terrazas et al., 2020; Voit et al., 2024). Thus, at around $M_{\rm*,c}$, all these coefficients are roughly constant, with only moderate redshift dependence introduced by the relation between $M_{\rm h}$ and $V_{\rm g}$. Consequently, the equilibrium equation predicts a power-law relation, $M_{\rm BH}\propto M_{*}^{5/3}$. At lower or higher mass scales, $f_{*}$ and $f_{\rm gas}$ cannot be viewed as a constant, and the power-law index may deviate from $5/3$. Note that the prediction of the scaling relation here is consistent with the semi-analytical model of Mo et al. (2024) and the observationally calibrated toy model of Hong et al. (2023), although they focus on the formation of bulge component and quenched galaxies, respectively.
@@
The example shown by the red dashed curve in Fig. 2 demonstrates the regulation over the equilibrium scaling relation. At $t_{\rm 23}$, the SMBH mass is slightly overshot relative to the power-law scaling with index $\gamma=5/3$. During Phase-3, the galaxy is pulled back towards the scaling, and finally comes to a region in the $_{\rm BH}-MM_{*}$ plane where other galaxies (shown by grey curves) converge. To statistically quantify the self-regulation nature, in Fig. 7, we show $k$, the power-law index of the $_{\rm BH}-MM_{*}$ relation during Phase-3 obtained from the best-fit function (Eq. 2), and $M_{\rm BH}(t_{23})/M_{*}(t_{23})$, the ratio of SMBH mass and stellar mass at the beginning of Phase-3, for individual galaxies, as well as the marginal distributions. The distribution of $M_{\rm BH}(t_{23})/M_{*}(t_{23})$ has a median of $\approx 10^{-2.4}$, approaching the typical value of galaxies at $z=0$. The distribution of $k$ has a median of $\sim 1.35$, smaller than the value of $5/3$ that Eq. (4.3.1) would imply if both $f_{\rm gas}$ and $f_{*}$ are constant. A strong anti-correlation between $k$ and $M_{\rm BH}(t_{23})/M_{*}(t_{23})$ is seen from the scattered points, and confirmed by the negative (rank-based) Spearman correlation coefficient. The dashed line in Fig. 7 shows a linear regression of $k$ on $\log\,\left[M_{\rm BH}(t_{23})/M_{*}(t_{23})\right]$, and has a slope of $\approx-3.0$. This slope indicates that, if a SMBH deviates from the equilibrium scaling relation (Eq. 4.3.1), it is pulled back once $M_{*}$ grows for about $0.3\,{\rm dex}$. For a typical main-sequence galaxy at $z\approx 0$ with sSFR $\sim 1\,{\rm Gyr}^{-1}$, this pull-back timescale is $\sim 2\,{\rm Gyr}$, shorter than the Phase-3 duration of most galaxies (Fig. 2). Hence, most SMBHs are well self-regulated during Phase-3, which erases the footprints of their early growth, including the seeding process. We have performed additional tests using star-forming galaxies at $z=0$ whose Phase-3 is well established, and found that the anti-correlation between $k$ and $M_{\rm BH}(t_{23})/M_{*}(t_{23})$ is still present, not limited by our sample selection (see §2). The self-regulation dominated growth also explains the tight relation between $M_{\rm BH}$ and $M_{*}$ at $z=0$, as shown by the black solid curve in Fig. 2. Observations also suggest a relation between $M_{\rm BH}$ and $M_{*}$ (or $M_{\rm*,bulge}$) in the local Universe (e.g. Kormendy & Ho, 2013; Greene et al., 2020; Zhang et al., 2024), which we will discuss in §5.1.
@@
The similar trend of sHAR, sSFR, and sBHAR during Phase-3 shown in the top left panel of Fig. 6 suggests that the rate of evolution of a galaxy in the $_{\rm BH}-MM_{\rm*}$ plane is controlled by the accretion rate of the host halo. The regulation by the AGN feedback bridges the growth of SMBH with its environment, transfers the growth pattern of the host halo to the galaxy and SMBH, and leads to the establishment of the scaling relation derived in Eq. (4.3.1). The sBHAR decreases with time during Phase-3, as it follows the sHAR that in general decreases with time. Once a certain mass threshold is reached, the mode of AGN feedback implemented in TNG makes a transition, as we will discuss in the following sections.
@@ 4.3.2 The trigger of the kinetic-mode feedback
Figure 8: Distribution of the properties relevant to the quenching of galaxies. Top panels show the distributions of individual galaxies in the $t_{\rm kin}$-$t_{\rm vmax,peak}$ (left) and $t_{\rm q}$-$t_{\rm vmax,\,peak}$ (right) planes, color-coded by their $M_{\rm h}$ at $z=0$. Bottom panels show the histograms of $t_{\rm q}-t_{\rm kin}$, the time lag of quenching after the trigger of the kinetic-mode feedback. The grey histogram in each panel shows the result for a subsample of galaxies in a given range of $M_{\rm h}$ at $z=0$. This subsample is further divided into three bins by $f^{[t_{\rm kin},t_{\rm q}]}_{\rm*,ex}$, the ex-situ growth of stellar mass between $t_{\rm kin}$ and $t_{\rm q}$, and their histograms are shown in different colors. The number of galaxies in each bin is labeled in the legend. See §4.3.2 and §4.3.3 for the definitions of these times and their implications for galaxy quenching.
@@
$\chi_{\rm thres}\equiv\min\left[0.002\left(\frac{M_{\rm BH}}{10^{8}\,{\rm M_{\odot}}}\right)^{2},0.1\right]\,$, (17)
@@
For low-mass SMBHs (a subsample with $M_{{\rm BH},z=0}\lesssim 10^{8}(2\times 10^{8})\,{\rm M_{\odot}}$, containing 38(578) galaxies, about 2.3 (40.5)% of our total sample, and with a median halo mass of $M_{{\rm h},z=0}=10^{11.8}(10^{12.2}){\rm M}_{\odot}$), the threshold $\chi_{\rm thres}$ is close to zero, and thus the transition occurs only if the sHAR approaches zero. A universal feature of halo assembly is the presence of two phases: a fast-accretion phase at high redshift, when the halo potential rapidly deepens, and a slow-accretion phase at low redshift, when the halo potential remains constant or even declines (e.g. Zhao et al., 2003). Thus, for the SMBH in a low-mass halo to trigger the kinetic-mode feedback, its host halo must transit to the slow phase. To verify that the halo transition indeed leads to the trigger of the kinetic-mode feedback, we follow Zhao et al. (2003) to define the halo transition time as $t_{\rm vmax,\,peak}$, the time when the maximum of the circular velocity, $V_{\rm max}$, reaches the peak, and we follow Weinberger et al. (2018, see their §3.1) to define whether the AGN feedback is in the thermal or kinetic mode. The last time when the AGN feedback switches from the thermal to kinetic mode is found, and denoted as $t_{\rm kin}$, for each galaxy in our sample. Only two galaxies in our sample are quenched at $z=0$ without activating the kinetic-mode feedback, and we remove them from the analyses involving $t_{\rm kin}$. We compare $t_{\rm kin}$ with $t_{\rm vmax,\,peak}$ for individual halos, and show the results in the top left panel of Fig. 8, color-coded according to their halo masses at $z=0$. For halos with $M_{{\rm h},z=0}\lesssim 10^{12.5}\,{\rm M_{\odot}}$, $t_{\rm vmax,\,peak}$ correlates well with $t_{\rm kin}$, indicating that the transition of the AGN feedback is indeed driven by the transition of halo accretion.
For massive SMBHs (a subsample with $M_{{\rm BH},z=0}\gtrsim 7\times 10^{8}\,{\rm M_{\odot}}$, containing 155 galaxies, about 10.9% of our total sample, and with a median halo mass of $M_{{\rm h},z=0}=10^{13.4}{\rm M}_{\odot}$), the threshold $\chi_{\rm thres}$ approaches the upper limit $0.1$. The transition of these halos can occur even when they are fast accreting. To derive the condition for their transition, we follow Dekel et al. (2023, see their §5.1) to approximate the specific accretion rate of a fast-accreting halo as
@@ 4.3.3 The quenching of galaxies
Figure 9: The sufficient condition for the quenching of galaxies. Left panel shows the distribution of galaxies in the $\dot{E}_{\rm cool,sub}$-$\dot{E}_{\rm kin}$ plane, where $\dot{E}_{\rm cool,sub}$ is the cooling rate of all non-star-forming gas cells in the subhalo, and $\dot{E}_{\rm kin}$ is the energy output rate from kinetic-mode AGN. Contours with a given color, from inner to outer, encompass 10% – 95% of the galaxies in our sample at a given time point ($t_{\rm q}$, $t_{\rm kin}$ or $z=0$) during their evolution histories along the main branch. Limited by the data storage, cooling rates are only available in the 20 “full snapshots” of TNG. Thus, the contour for $t_{\rm kin}$ (or $t_{\rm q}$) only includes the galaxies whose $t_{\rm kin}$ ($t_{\rm q}$) is in the “full snapshots”, which reduces the sample size to 199 (304). Dashed line indicates where these two rates are equal. Right panel shows the histograms of the ratio, $\dot{E}_{\rm kin}/\dot{E}_{\rm cool,sub}$, at different time points. Dashed lines indicate the medians of the distributions. See §4.3.3 for the details.
The feedback energy in the kinetic mode implemented by TNG affects the gas effectively by both ejection and prevention, and can quickly reduce the amount of star-forming gas in the galaxy and beyond (e.g. Zinger et al., 2020; Shi et al., 2022; Martín-Navarro et al., 2021). Consequently, galaxies are expected to be quenched quickly after the transition of the feedback mode. In the top right panel of Fig. 8, we show $t_{\rm q}$, the time when a galaxy is quenched (see §2 for the definition), and compare it to $t_{\rm vmax,\,peak}$ for individual galaxies. For all galaxies, $t_{\rm q}$ is systematically later (larger) than $t_{\rm kin}$, indicating that the kinetic-mode AGN feedback is a necessary condition for quenching central galaxies, and that galaxies may sustain their star formation for some time after the trigger of the kinetic-mode feedback.
@@
where $\Delta M_{\rm*,ex}^{[t_{\rm kin},t_{\rm q}]}$ is the mass of stars formed ex-situ during the time interval between $t_{\rm kin}$ and $t_{\rm q}$ and fallen into the galaxy before $t_{\rm q}$ via either stripping or merger, and $M_{*}(t_{\rm q})$ is the stellar mass of the galaxy at $t_{\rm q}$. We utilize the published catalog generated by Rodriguez-Gomez et al. (2015) and Rodriguez-Gomez et al. (2016) to identify the ex-situ stars for each galaxy, defined as those bound to the galaxy but formed outside the main branch of the subhalo merger tree rooted in the galaxy. The distribution of the time lag of quenching for each subsample is shown in the bottom panels of Fig. 8 by a colored histogram. For galaxies with $f_{\rm*,ex}^{[t_{\rm kin},t_{\rm q}]}<5\%$, the time lag is sharply peaked within $1\,{\rm Gyr}$, regardless of the halo mass. This demonstrates that the kinetic-mode feedback is indeed effective in quenching the galaxy in the absence of merger-induced gas inflow. However, for galaxies with strong mergers ($f_{\rm*,ex}^{[t_{\rm kin},t_{\rm q}]}>30\%$), star formation can sustain a long time, even if the kinetic-mode feedback is triggered. The quenching of a galaxy is thus delayed to a time when the frequency of mergers falls below a certain level, depending on the sSFR threshold used to define the quenching (§2). A similar conclusion was reached by Terrazas et al. (2020), where the cumulative energy of AGN is computed and compared to the binding energy of the gas in a galaxy. Galaxies are found to be quenched only when the feedback energy exceeds about 100 times the binding energy of the gas, indicating that most of the feedback energy is lost due to radiative cooling.
@@
The cease of the merger marks the end of the stochastic creation of star-forming gas, leaving the cooling of hotter gas in a halo being the only channel for the central galaxy to maintain its star formation. In this case, the gas surrounding the SMBH is dominated by that in the hot phase, as shown in the bottom right panel of Fig. 6. If the energy injection by the kinetic-mode feedback remains strong even when the amount of gas available for the SMBH to accrete is small, as is the case in TNG because of its large feedback efficiency at low accretion rates (Weinberger et al., 2017, see their §2.3), the amount of the gas within the galaxy will continue to decrease due to both ejection and prevention, causing the rates of gas cooling and star formation to slow down. As shown in the top-right panel of Fig. 6, the effective sound speed of the gas during Phase-3 continues to increase, while the density decreases, both leading to a decrease in $A_{\rm s}$ and thus to the growth of the SMBH and the strength of its feedback. Such an evolution stops once the feedback of AGN is reduced to a level comparable to the cooling rate of the gas. This final stage, where cooling, star formation, SMBH growth, and AGN feedback are all slow, produces a quenched state that is consistent with observations.
@@
The above argument leads to an equality that characterizes the condition for the kinetic-mode AGN feedback to maintain the quenching of a galaxy without significant mergers, which is that the feedback balances the cooling of gas:
@@
where $\dot{E}_{\rm cool,sub}$ is the total cooling rate of all non-star-forming gas cells within the host subhalo of the galaxy, $\dot{E}_{\rm kin}\equiv\eta_{\rm kin}\dot{M}_{\rm BH}c^{2}$ is the energy output rate of the kinetic-mode AGN, and $\eta_{\rm kin}={\rm min}\left[\rho/(0.05\rho_{\rm SF,\,thresh}),\,0.2\right]$ is a coupling efficiency (Weinberger et al., 2018, see their Eq. 5). The inclusion of the subhalo gas outside the galaxy is necessary for TNG, as its kinetic-mode feedback works through both ejection and prevention over a large scale. The contours in the left panel of Fig. 9 show the distribution of galaxies in the $\dot{E}_{\rm cool,sub}$-$\dot{E}_{\rm kin}$ plane at different epochs, while the histograms in the right panel show the marginal distributions of $\dot{E}_{\rm kin}/\dot{E}_{\rm cool,sub}$, both in logarithmic scales. As expected, at the trigger time of the kinetic-mode feedback, $t=t_{\rm kin}$, both $\dot{E}_{\rm kin}$ and $\dot{E}_{\rm cool,sub}$ are systematically larger than those at later times. In particular, $\dot{E}_{\rm kin}$ is larger than $\dot{E}_{\rm cool,sub}$, indicating a non-equilibrium and the continuing reduction of gas. At the quenching point, $t=t_{\rm q}$, the two rates become balanced, both significantly smaller than those at $t_{\rm kin}$, implying a stable state of quenching and the transition of the galaxy into the final phase.